Suppose that the demand for labor is given by P equals 30 minus Q and the supply of labor is given by P equals 2Q. Suppose the government decides to impose a wage of P equals 26. Which of the following statements best describes the outcome of such a price?

In comparison with the equilibrium outcome, an imposed price of 26 results in an unemployment level of 13 and a deadweight loss of 150.

In comparison with the equilibrium outcome, an imposed price of 26 results in an unemployment level of 13 and a deadweight loss of 54.

In comparison with the equilibrium outcome, an imposed price of 26 results in an unemployment level of 9 and a deadweight loss of 54.

In comparison with the equilibrium outcome, an imposed price of 26 results in an unemployment level of 9 and a deadweight loss of 150.
Answer #1

Demand for labor: P=30-Q

Supply of labor: P=2Q

Equilibrium arises where:

Demand = Supply

30-Q= 2Q

30= 3Q

Q*= 10 Equilibrium quantity of labor

P*= 2(10)= 20 Equilibrium wage

Suppose the government decides to impose a wage of P equals 26(Pg=26):

Demand for labor: P= 30-Q

26 = 30-Q

Qd= 30-26= 4 Demand for labor

Supply of labor: P=2Q

26=2Q

Q= 13 Supply of labor

Unemployment= Supply - demand= 13-4= 9

When demand for labor is 4. At this quantity of labor, supplier ready to receive: Ps=2(4)= 8

Deadweight loss= (1/2)(Pg-Ps)(Q*-Qd)= (1/2)(26-8)(10-4)= 9 x 6= 54

Wage(P) Supply Deadweight loss 26 E 20 8 Demand 4 10 13 Quantity Unemployment

In comparison with the equilibrium outcome, an imposed price of 26 results in an unemployment level of 9 and a deadweight loss of 54.